# Anyview: Generalizable Indoor 3D Object Detection with Variable Frames

### v1

@@ I Introduction
3D object detection is a fundamental scene understanding problem for robotic manipulation [45, 57], autonomous driving [46], and AR/VR [54, 27], which aims to detect the 3D bounding boxes and semantic labels from point clouds or images. Due to the different types of sensors used in different application scenarios, 3D object detection methods usually vary a lot for indoor [33, 13] and outdoor [61] scenes. We focus on indoor 3D object detection, where the mainstream sensor is RGB-D cameras and the scenes are crowded with objects of multiple categories and sizes.
@@
Although great improvement in performance has been achieved by advanced architecture design, existing methods train and evaluate 3D detectors on only fixed frames of input data, such as monocular RGB-D images (on SUN-RGBD [37] benchmark) and scene-level reconstructed point clouds (sampled from meshes) from multi-view RGB-D images (on ScanNet [5] benchmark). There is still a huge gap between these benchmarks and practical applications. Specifically, due to the various task budgets in practical application scenes, the number of frames input to the 3D detectors is dynamic, and previous frameworks trained on fixed frame inputs are challenging to generalize. The cost of training diverse models for tasks with various sizes of inputs is huge, severely limiting the deployment of the models on edge devices. In order to solve this problem, it is necessary to further research and develop more efficient generalized indoor 3D object detection with variable frames. For example, in online tasks such as robot navigation [12, 49, 18], the information captured by the agent is strictly limited by a cost budget constraint, which leads to a varying number of input frames. As we will show in Section III-A, previous models generalize poorly across different input scales, which brings a huge burden as we need to prepare a series of models trained on different scales of input data to cope with various scenarios.
@@
In practical scenes, the raw input to the indoor 3D object detectors is the RGB-D images with variable frame numbers. The detectors require the ability to process inputs with various frame numbers at fixed parameters to complete different tasks. To simulate this, we propose a novel processing framework named AnyView for indoor 3D object detection as shown in Fig. 1. Previous approaches require a complete reconstruction of the scene point cloud through dense RGB-D images before input. This limitation results in the models only being able to process a fixed frame number of RGB-D image inputs, leading to weak generalization. In addition, a large amount of resources are wasted in the scene reconstruction process. Compared with previous approaches, our proposed AnyView can handle variable frame numbers of input data to satisfy the cost budgets of various practical applications. Specifically, we first unify the modality of input data as sparse multi-view RGB-D images and the corresponding camera parameters, which are compatible with current benchmark datasets (e.g. ScanNet [5] and SUN-RGBD [37]). Instead of concatenating multi-view point clouds into a whole and extracting a fixed number of global features of scenes, we learn geometric clues for each frame independently and design a transformer-based architecture to merge and refine the extracted features of various frames to achieve the ability to efficiently process variable frame inputs. Specifically, we propose a geometric learner to mine local geometric structural features and employ a spatial mixtures module to fuse local details with global semantics. In this way, AnyView is able to extract rich representations for input scenes consisting of any number of frames, bridging the gap between the local geometric features of each frame and the global semantic features. We further propose the randomized view and rectangular dropout data enhancement strategies to encourage the model to adapt to various scales of inputs and ignore frames with less geometric information, respectively. In addition, to further enhance the generalization of AnyView and consistent point cloud feature density, we propose a dynamic token strategy to adaptively adjust the number of extracted features per frame according to input frame count, which ensures the parameter compatibility and uniformity of the point cloud feature scale. AnyView fuses the multiview point cloud with camera parameters and refines the initial proposal with a transformer decoder based on the extracted scene representations. Extensive experiments on ScanNet show that our method outperforms previous methods in both accuracy and generalizability with a similar amount of parameters. In the online detection simulation on the ScanNet dataset, AnyView achieves 55.23%mAP@25, which is 14.14% higher than 3DETR [23].
@@ II Related Work
Advances in 3D sensors have led to a surge of methods designed for 3D object detection with various modality. The large amount of related works can be divided into three groups based on the modality of 3D data. We also briefly discuss the sensors and datasets corresponding to these modalities.
@@
Point-based 3D object detection: This kind of methods take in 3D scenes represented by pure point clouds, which are acquired by LIDAR sensor [10] or 3D reconstruction from multi-view RGB-D images [5, 60, 9]. Early 3D object detection methods mainly include sliding-window methods [38] and template-based methods [17, 19]. Deep learning-based end-to-end 3D object detection methods began to emerge in recent years, which are mainly based on PointNet [31, 32] or sparse CNN [11] backbones. PointNet-based methods [35, 29, 23, 47] consume point clouds directly with the set abstraction operation [32], which enables flexible receptive fields for 3D feature learning. Sparse CNN-based methods [60, 42, 55, 8] project the point clouds to regular grids to be processed by advanced 2D or 3D CNN architectures. Tang et al. [41] further combine the two representations to learn more discriminative features. Liu et al. [20] proposed a dynamic point cloud joint detection framework, which uses spatio-temporal information to improve the problem of target loss due to severe occlusions.
Recently numerous works have employed bird’s eye view BEV to improve the point cloud feature extraction backbone network. Deng et al. [6] proposed the use of point cloud projection onto perspective views and BEV to capture multiview features, which employed a novel bilaterally guided multiview fusion block to phantom 3D representations leveraging the complementary information in perspective views and BEV. An et al. [1] designed a saliency detection task to enhance the detector for sparse point cloud object recognition, exploiting voxels and BEV saliency maps as asymptotic attention to resist redundant features in the background region. Xie et al. [51] proposed a novel local global feature adaptive aggregation mechanism to fuse contextual information at a fine-grained level and improve the quality of proposals with weighted relational awareness.
Several researchers have enhanced point cloud feature extraction by fusing information from multiple sources. Liu et al. [20] proposed a dynamic point cloud joint detection framework, which uses spatio-temporal information to improve the problem of target loss due to severe occlusions. Shan et al. [15] proposed to utilize the PTT module for modeling feature interactions between the voting phase and proposal generation phase of object detection, which enhances the feature extraction of sparse point cloud objects. However, point-based detection models have high inference latency due to inefficient sampling algorithms.
@@
RGBD-based 3D object detection: RGBD camera is the mainstream 3D sensor for indoor scene understanding tasks. Existing RGBD-based 3D object detection methods take in monocular RGB-D image [24, 37] or point clouds reconstructed from multi-view RGB-D images (mentioned above). Here we only review the former methods which make full use of both RGB and depth information. Prior methods broadly fall into three categories: 2D-driven, 3D-driven, and modal fusion. 2D-driven methods [7, 52] first detect object in images and then use the results to assist search in 3D space. Pahwa et al. [25] proposed to utilize per-frame depth information and multi-view scene information to obtain accurate 3D object suggestions and generate 3D bounding boxes for potential 3D regions of interest. Kundu et al. [16] presented a microscopic rendering and comparison loss by employing voxel patterns to represent 3D objects, which allows 3D shapes and poses to be learned via 2D supervision.
The most common approach to 3D-driven is to feed image information into a 3D feature extraction network as an additional channel to a point cloud or voxel. Song et al. [38] localized objects on a voxelized point cloud by sliding a 3D detection window, and category information was obtained from RGB pixels attached to the point cloud. Song et al. [39] proposed to process RGB-D inputs by sparse convolution and generate 3D proposals, and subsequently determine object categories by a designed joint object recognition network. However, this type of simple point cloud and image combination approach may destroy the fine-grained local geometric structural features of the point cloud, resulting in inefficient fusion between the point cloud and image modalities.
Modal fusion methods [22, 28] use the geometric correspondence between 2D and 3D to guide the predictions of the network and enrich the representation by concatenating multi-modal features. Zhu et al. [62] employed virtual point cloud images to effectively bridge the resolution gap between image modality and point cloud modality to address the information loss due to multimodal data fusion. Bai et al. [2] employed Transformer to adaptively fuse 3D point cloud proposals with image features and processed difficult samples in the point cloud with an image-guided strategy, effectively improving the robustness of low-quality images. However, there is still a huge performance gap between RGBD-based methods and point cloud-based methods.
Multi-sensor 3D object detection: Recently, multi-sensor fusion has aroused increased interest in the 3D detection community. Depending on the fine-grained level of the fusion approach, existing approaches can be classified into region-level and point-level fusion methods.
Region-level fusion methods generate object proposals in 3D space and project them to images [4] or lift image proposals into a 3D frustum to reduce the search space in 3D space [30, 36]. Qi et al. ] [30] implemented multimodal inference directly employing a 2D object detector with a 3D point cloud feature extraction network, this framework was able to maintain high accuracy even in the presence of strong occlusions or very sparse points and inspired a series of subsequent work Wang et al. [48] proposed to project region proposals on 2D images as frustums in 3D space and utilize 3D backbone network extraction to achieve region-level feature fusion. Liu et al. [21] proposed a fusion module based on a depth-attention mechanism to achieve multimodal feature differential fusion at various depths and addressed the unstable efficiency of modal fusion of point clouds and images. Paigwar et al. [26] extended the fusion approach to pillar representation and proposed a novel masked point cloud approach to improve object localization accuracy.
Point-level fusion methods, on the other hand, usually enhance the LIDAR points with the image semantic features. There are several researches aiming to enhance the input point cloud with image features and then employ a 3D object detector to process the enhanced point cloud for higher accuracy. The image features can be fused to input LIDAR points [44, 56] or middle feature points [59, 50]. Xu et al. [53] utilized point-to-pixel projection matrices to attach 2D semantic segmentation labels to 3D point clouds and employed an attentional mechanism to adaptively fuse semantic information to significantly improve model performance. Yin et al. [56] proposed to generate a dense 3D virtual point cloud via a 2D images object detector to solve the point cloud sparsity problem and this approach can be naturally integrated into any LiDAR data. Vora et al. [44] proposed to project 3D point cloud information to the 2D image semantic segmentation results and can feed the integrated information back to any 3D point cloud object detection models, which provides a better trade-off between accuracy and speed. After that, point-based methods are adopted to detect objects from the decorated point clouds. However, effectively fusing information from multiple modalities remains an open challenge.
@@ III Analysis
In this section, we first describe the problem statement from existing benchmarks for indoor 3D object detection. Then we propose a more unified and reasonable setting for indoor 3D object detection task. Finally we show the failing case of previous models in practical scenarios.
@@ IV Approach
Figure 2: The framework of AnyView. Given multi-view RGB-D images of variable number, we first convert the depth maps to single view point clouds according to intrinsic camera parameters. Then a shared geometry learner is applied to extract scene proxies for each view, which are local feature descriptors independent of input scale. Scene proxies from different views are mixed into world coordinates according to the extrinsic camera parameters. We adopt transformer encoder to refine the scene proxies while keeping their input scale-independent property by self-attention mechanism. A transformer decoder is used to refine the object proposals, where the initial proposals are sampled from the whole scene by FPS. The dashed blue box indicates the number of inside elements is changable during inference.
In this section, we propose a new framework called AnyView for detecting 3D objects from multi-view RGB-D images. We first show the overall framework of AnyView. Then we introduce the input scale-independent training strategy. Finally we detail the dynamic token technique for accuracy-computation tradeoff during inference time.
@@ IV-A Overall Framework
Interactions among scene proxies: We obtain richer scene representations through interactions among scene proxies. In order to keep the input scale-independent property of scene proxies, we hope the interaction to be linear combination $\mathcal{P}_{ij}=\sum_{i=1}^{N}\sum_{j=1}^{T}{\alpha_{ij}\mathcal{P}_{ij}}$, where $\sum_{i}\sum_{j}{\alpha_{ij}}=1$. In addition, since the number of view ($N$) is variable, the number of interacting scene proxies is also variable. Benefiting from the nature of self-attention, these requirements can be elegantly achieved by transformer [43].
@@
where $\mathcal{M}$ represents the MLP and $\mathcal{F}$ represents the farthest point sampling algorithm. The refined queries are converted into box parameters and supervised following the procedure in 3DETR [23].
@@ IV-B Scale-independent Training
Although the architecture design of AnyView can extract features independent of input scale and detect 3D objects from any number of views, we find proper data augmentation is essential for training the model, which are listed below.
@@
Global random cuboid: The geometry structure of single view point clouds may vary from view to view. For example, when the distance between the RGB-D camera and the scene varies, the density of point clouds varies as well. When one view only contains floors and walls, the scene proxies extracted from it are uninformative. To allow the network to handle depth maps taken from various shooting situations and learn to ignore scene proxies with less information, we randomly crop a cuboid [58] from the fused point clouds $P={\rm\mathcal{C}}(R_{i}^{c}\cdot P_{ci}+T_{i}^{c})$ and set other points outside the cuboid to $(0,0,0)$. After that, we transform the point clouds back to the camera coordinates of each view and keep $(0,0,0)$ unchanged. This strategy mak some views lose geometric information but still provide valid scene proxies, which effectively reduces overfitting and improves the generalizability of AnyView.
@@ IV-C Dynamic Tokens for Inference
AnyView extracts $N_{T}$ scene proxies for a $N$-view input scene. In our setting, $N$ may vary from 1 to 50 during inference time, which results in an order of magnitude change in the number of scene proxies. In previous work such as VoteNet [29] and 3DETR [23], the encoder of detector extract 1024 or 2048 seed points for one scene, no matter monocular RGB-D input or scene-level point clouds input. To reach this order of magnitude, $T$ should be around 40 when $N=50$ and be around 2000 when $N=1$. For the former case, AnyView only extracts 40 scene proxies for monocular input, which is not enough to fully represent the scene. While for the latter case, the number of scene proxies will be larger than 10000 when there are more than 10 views. This heavy computation cost is unaffordable in practical applications.
@@ V Experiment
In this section, we conduct experiments to show the generalizability of our approach. We first describe the datasets and experimental settings. Then we comprehensively compare our approach with previous indoor 3D object detection methods on a series of benchmarks. Visualization results are provided for more intuitive comparison. Finally we design several ablation studies to verify the effectiveness of our architecture design and training/inference strategies.
@@
ImVoteNet
ImVoxelNet
3DETR-m
AnyView
1.8M
104.6M
7.4M
7.6M
@@ V-A Experiments Setup
Datasets and benchmarks: We conduct experiments on the ScanNet [5] dataset, which is a richly annotated dataset of indoor scenes with 1201 training scenes and 312 validation scenes. For each scene, ScanNet provides RGB-D video as well as the reconstructed mesh. We uniformly sample 50 frames from the video for each scene, which serves as the maximum data available for training and evaluating in our proposed setting. Meanwhile, we further validate that the proposed approach has the ability to accurately detect objects in practical deployment scenes by verifying the performance of each model on the online detection task. We set three benchmarks as defined in Section III-A. ScanNet-Rec is the mainstream scene-level benchmark used by previous works [14, 29, 23], whose input data is reconstructed point cloud of a whole scene and ground-truth is generated by computing the axis-aligned bounding boxes [14] for objects in 18 selected categories. In addition, we designed ScanNet-Online benchmarks to explore the performance of previous 3D object detection models with AnyView in practical deployment scenes. Specifically, the input to ScanNet-Online is each frame in a sequence of RGB-D images, and the model needs to dynamically detect objects in the scene based on the new input frames. ScanNet-MV shares the same ground truth with ScanNet-Rec, while the input data are multi-view RGB-D images. If not additionally mentioned, the default number of views is 50. ScanNet-MV is also a scene-level benchmark, which means all RGB-D images from one scene are considered as a single input sample for the network. ScanNet-SV is a monocular RGB-D benchmark with a single view RGB-D image as input data. For each image, we select the bounding boxes whose center points are within the image from the corresponding scene as its ground truth.. ScanNet-Online utilizes the ScanNet-MV benchmark in the 50 views setting, and the input is each frame of sequential RGB-D images.
@@
TABLE III: 3D object detection results (mAP@0.25 and mAP@0.5) and corresponding training/inference setting of different methods on scene-level and monocular benchmarks. Contents in bracket indicate the modality used by the methods.
@@
TABLE IV: 3D object detection results (mAP@0.25 and mAP@0.5) of different models on ScanNet-SV and ScanNet-MV. Here ScanNet-MV is further divided according to the number of views and how these views are combined. Uniform/Continuous means the views are sampled uniformly/adjacently from the whole 50 views. Gray cells show on which benchmark the model is trained.
@@
 5
10
15
30
40
50
Uniform ImVoteNet 31.6 / 13.5
30.2 / 13.7
37.3 / 17.9
40.3 / 18.5
43.5 / 20.6
42.9 / 20.4
43.3 / 20.4
3DETR-m(10)
29.8 / 12.5
39.7 / 20.0
49.1 / 28.0
50.8 / 29.0
50.3 / 28.9
49.6 / 28.9
48.9 / 27.7
3DETR-m(30)
29.1 / 13.1
36.8 / 19.5
50.1 / 29.6
53.1 / 31.4
55.6 / 33.9
55.7 / 34.3
55.6 / 34.7
3DETR-m(50)
27.7 / 12.8
34.9 / 17.9
48.6 / 29.0
52.5 / 33.3
55.3 / 36.3
55.6 / 35.9
54.7 / 35.3
AnyView(10)
31.3 / 13.7
42.2 / 20.4
51.3 / 27.2
52.8 / 29.0
53.2 / 29.9
54.9 / 30.4
54.8 / 30.1
AnyView(30)
30.5 / 12.7
39.8 / 20.1
52.4 / 29.2
57.1 / 31.6
59.1 / 35.2
59.9 / 36.8
60.1 / 37.4
AnyView(50)
32.7 / 15.3
45.0 / 22.0
53.8 / 30.5
57.5 / 33.5
59.5 / 35.7
60.6 / 35.7
60.7 / 35.8
Continuous ImVoteNet 31.6 / 13.5
18.1 / 8.8
25.2 / 13.5
30.4 / 15.7
40.5 / 18.4
43.1 / 20.0
43.3 / 20.4
3DETR-m(10)
29.8 / 12.5
19.9 / 9.9
27.7 / 15.3
34.0 / 18.1
44.2 / 24.1
48.7 / 27.8
48.9 / 27.7
3DETR-m(30)
29.1 / 13.1
18.5 / 9.6
26.6 / 15.1
35.4 / 19.7
48.0 / 29.4
54.0 / 33.6
55.6 / 34.7
3DETR-m(50)
27.7 / 12.8
17.9 / 9.7
26.6 / 15.8
34.9 / 19.8
46.9 / 28.8
52.7 / 33.0
54.7 / 35.3
AnyView(10)
31.3 / 13.7
20.6 / 10.6
30.2 / 15.7
36.0 / 17.7
47.9 / 26.1
53.6 / 28.7
54.8 / 30.1
AnyView(30)
30.5 / 12.7
19.9 / 10.1
29.9 / 16.5
36.8 / 21.4
52.3 / 30.0
57.1 / 34.3
60.1 / 37.4
AnyView(50)
32.7 / 15.3
22.1 / 11.3
32.9 / 18.6
40.6 / 21.8
54.4 / 29.8
58.7 / 34.4
60.7 / 35.8
Figure 3: Visual results on ScanNet. We compare the predictions of 3DETR-m and AnyView after NMS with the ground-truth bounding boxes on different scales of input. Top: the whole scene consisting of 50 views, Bottom: a part of scene consisting of 10 adjacent views. AnyView successfully detects all the doors without color information, which shows its strong ability for extracting local geometric information. The predictions of AnyView is also more consistent across different input scales.

### v2

# Anyview: General Indoor 3D Object Detection with Variable Frames

@@ I INTRODUCTION
3D object detection is a fundamental scene understanding problem for robot manipulation [19], navigation [25], which aims to detect the 3D bounding boxes and semantic labels from point clouds or images. Due to the different types of sensors used in different application scenarios, 3D object detection methods usually vary a lot for indoor [23] and outdoor [20] scenes. We focus on indoor 3D object detection, where the mainstream sensor is RGB-D cameras and the scenes are crowded with objects of multiple categories and sizes.
@@
Although great improvement in performance has been achieved by advanced architecture design, existing methods train and evaluate 3D detectors on only fixed frames of input data, such as monocular RGB-D images (on SUN-RGBD [14] benchmark) and scene-level reconstructed point clouds (sampled from meshes) from multi-view RGB-D images (on ScanNet [2] benchmark). There is still a huge gap between these benchmarks and practical applications. Specifically, due to the various task budgets in practical application scenes, the number of frames input to the 3D detectors is dynamic, and previous frameworks trained on fixed frame inputs are challenging to generalize. The cost of training diverse models for tasks with various sizes of inputs is huge, severely limiting the deployment of the models on edge devices. In order to solve this problem, it is necessary to further research and develop more efficient generalized indoor 3D object detection with variable frames. For example, in online tasks such as robot navigation [25], the information captured by the agent is strictly limited by a cost budget constraint, which leads to a varying number of input frames. As we will show in Section III-A, previous models generalize poorly across different input scales, which brings a huge burden as we need to prepare a series of models trained on different scales of input data to cope with various scenarios.
@@
In practical scenes, the raw input to the indoor 3D object detectors is the RGB-D images with variable frame numbers. The detectors require the ability to process inputs with various frame numbers at fixed parameters to complete different tasks. To simulate this, we propose a novel processing framework named AnyView for indoor 3D object detection as shown in Fig. 1. Previous methods require complete scene reconstruction from dense RGB-D images, limiting the model to a fixed number of inputs, which weakens generalization and wastes resources. In contrast, AnyView handles variable frame numbers, making it adaptable to different application budgets. We unify input data as sparse multi-view RGB-D images with camera parameters, compatible with datasets like ScanNet [2] and SUN-RGBD [14]. Instead of concatenating point clouds and extracting fixed global features, AnyView learns geometric clues per frame and uses a transformer-based architecture to efficiently merge and refine these features, enabling processing of variable frame inputs. We propose a geometric learner to extract local geometric features and a spatial mixtures module to combine these with global semantics, which enables AnyView to generate rich representations from scenes with any number of frames, bridging local geometric and global semantic features. We introduce randomized view and rectangular dropout strategies to help the model adapt to varying input scales and discard frames with less geometric information. Additionally, a dynamic token strategy ensures consistent feature density and parameter compatibility across point clouds. Extensive experiments on ScanNet show that our method outperforms previous methods in both accuracy and generalizability with a similar amount of parameters.
@@ II Related Work
Point-based 3D object detection: This kind of methods take in 3D scenes represented by pure point clouds, which are acquired by LIDAR sensor or 3D reconstruction from multi-view RGB-D images [13]. Early 3D object detection methods mainly include sliding-window methods [15] and template-based methods [6]. Deep learning-based end-to-end 3D object detection methods began to emerge in recent years, which are mainly based on PointNet [10] or sparse CNN [24, 4] backbones. PointNet-based methods [18] consume point clouds directly with the set abstraction operation [11], which enables flexible receptive fields for 3D feature learning. Sparse CNN-based methods [17] project the point clouds to regular grids to be processed by advanced 2D or 3D CNN architectures. However, point-based detection models have high inference latency due to inefficient sampling algorithms.
@@
RGBD-based 3D object detection: RGBD camera is the mainstream 3D sensor for indoor scene understanding tasks. Existing RGBD-based 3D object detection methods take in monocular RGB-D image [22] or point clouds reconstructed from multi-view RGB-D images (mentioned above). Here we only review the former methods which make full use of both RGB and depth information. Prior methods broadly fall into three categories: 2D-driven, 3D-driven, and modal fusion. 2D-driven methods [21] first detect object in images and then use the results to assist search in 3D space. Kundu et al. [5] presented a microscopic rendering and comparison loss by employing voxel patterns to represent 3D objects, which allows 3D shapes and poses to be learned via 2D supervision. However, there is still a huge performance gap between RGBD-based methods and point cloud-based methods. The most common approach to 3D-driven is to feed image information into a 3D feature extraction network as an additional channel to a point cloud or voxel. Song et al. [15] localized objects on a voxelized point cloud by sliding a 3D detection window, and category information was obtained from RGB pixels attached to the point cloud. However, this type of simple point cloud and image combination approach may destroy the fine-grained local geometric structural features of the point cloud, resulting in inefficient fusion between the point cloud and image modalities.
@@ IV Approach
Figure 2: The framework of AnyView. Given multi-view RGB-D images of variable number, we first convert the depth maps to single view point clouds according to intrinsic camera parameters. Then a shared geometry learner is applied to extract scene proxies for each view, which are local feature descriptors independent of input scale. Scene proxies from different views are mixed into world coordinates according to the extrinsic camera parameters. We adopt transformer encoder to refine the scene proxies while keeping their input scale-independent property by self-attention mechanism. A transformer decoder is used to refine the object proposals, where the initial proposals are sampled from the whole scene by FPS. The dashed blue box indicates the number of inside elements is changeable during inference.
@@ IV-A Overall Framework
Interactions among scene proxies: We obtain richer scene representations through interactions among scene proxies. In order to keep the input scale-independent property of scene proxies, we hope the interaction to be linear combination $\mathcal{P}_{ij}=\sum_{i=1}^{N}\sum_{j=1}^{T}{\alpha_{ij}\mathcal{P}_{ij}}$, where $\sum_{i}\sum_{j}{\alpha_{ij}}=1$. In addition, since the number of view ($N$) is variable, the number of interacting scene proxies is also variable. Benefiting from the nature of self-attention, these requirements can be elegantly achieved by transformer.
@@
where $\mathcal{M}$ represents the MLP and $\mathcal{F}$ represents the farthest point sampling algorithm. The refined queries are converted into box parameters and supervised following the procedure in 3DETR.
@@ IV-B Scale-independent Training
Global random cuboid: The geometry structure of single view point clouds may vary from view to view. For example, when the distance between the RGB-D camera and the scene varies, the density of point clouds varies as well. When one view only contains floors and walls, the scene proxies extracted from it are uninformative. To allow the network to handle depth maps taken from various shooting situations and learn to ignore scene proxies with less information, we randomly crop a cuboid from the fused point clouds $P={\rm\mathcal{C}}(R_{i}^{c}\cdot P_{ci}+T_{i}^{c})$ and set other points outside the cuboid to $(0,0,0)$. After that, we transform the point clouds back to the camera coordinates of each view and keep $(0,0,0)$ unchanged. This strategy makes some views lose geometric information but still provide valid scene proxies, which effectively reduces overfitting and improves the generalizability of AnyView.
@@ IV-C Dynamic Tokens for Inference
AnyView extracts $N_{T}$ scene proxies for a $N$-view input scene. In our setting, $N$ may vary from 1 to 50 during inference time, which results in an order of magnitude change in the number of scene proxies. In previous work such as VoteNet [9] and 3DETR, the encoder of detector extract 1024 or 2048 seed points for one scene, no matter monocular RGB-D input or scene-level point clouds input. To reach this order of magnitude, $T$ should be around 40 when $N=50$ and be around 2000 when $N=1$. For the former case, AnyView only extracts 40 scene proxies for monocular input, which is not enough to fully represent the scene. While for the latter case, the number of scene proxies will be larger than 10000 when there are more than 10 views. This heavy computation cost is unaffordable in practical applications.
@@ V Experiment
ImVoteNet ImVoxelNet 3DETR-m AnyView
1.8M 104.6M 7.4M 7.6M
@@ V-A Experiments Setup
Datasets and benchmarks: We conduct experiments on the ScanNet [2] dataset, which is a richly annotated dataset of indoor scenes with 1201 training scenes and 312 validation scenes. For each scene, ScanNet provides RGB-D video as well as the reconstructed mesh. We uniformly sample 50 frames from the video for each scene, which serves as the maximum data available for training and evaluating in our proposed setting. Meanwhile, we further validate that the proposed approach has the ability to accurately detect objects in practical deployment scenes by verifying the performance of each model on the online detection task. We set three benchmarks as defined in Section III-A. ScanNet-Rec is the mainstream scene-level benchmark used by previous works [3, 9, 7], whose input data is reconstructed point cloud of a whole scene and ground-truth is generated by computing the axis-aligned bounding boxes [3] for objects in 18 selected categories. In addition, we designed ScanNet-Online benchmarks to explore the performance of previous 3D object detection models with AnyView in practical deployment scenes. Specifically, the input to ScanNet-MV shares the same ground truth with ScanNet-Rec, while the input data are multi-view RGB-D images. If not additionally mentioned, the default number of views is 50. ScanNet-MV is also a scene-level benchmark, which means all RGB-D images from one scene are considered as a single input sample for the network. ScanNet-Online is each frame in a sequence of RGB-D images, and the model needs to dynamically detect objects in the scene based on the new input frames. Compared with ScanNet-MV, ScanNet-Online inputs each frame of RGBD images in sequence rather than directly inputting all frames at the scene level, which is more suitable for practical deployment requirements. ScanNet-SV is a monocular RGB-D benchmark with a single view RGB-D image as input data. For each image, we select the bounding boxes whose center points are within the image from the corresponding scene as its ground truth.
@@
TABLE III: 3D object detection results (mAP@0.25 and mAP@0.5) on scene-level and monocular benchmarks. Contents in brackets indicate the method modality, where (D) denotes depth only and (RGBD) means added RGB images.
@@
TABLE IV: 3D object detection results (mAP@0.25/0.5) on ScanNet-SV/ScanNet-MV. Uniform/Continuous means the views are sampled uniformly/adjacent from the whole 50 views. Gray cells show on which benchmark the model is trained. Contents in brackets represent training views number.
@@
 5 10 15 30 40 50
Uniform ImVoteNet 31.6 / 13.5 30.2 / 13.7 37.3 / 17.9 40.3 / 18.5 43.5 / 20.6 42.9 / 20.4 43.3 / 20.4
3DETR-m(10) 29.8 / 12.5 39.7 / 20.0 49.1 / 28.0 50.8 / 29.0 50.3 / 28.9 49.6 / 28.9 48.9 / 27.7
3DETR-m(30) 29.1 / 13.1 36.8 / 19.5 50.1 / 29.6 53.1 / 31.4 55.6 / 33.9 55.7 / 34.3 55.6 / 34.7
3DETR-m(50) 27.7 / 12.8 34.9 / 17.9 48.6 / 29.0 52.5 / 33.3 55.3 / 36.3 55.6 / 35.9 54.7 / 35.3
AnyView(10) 31.3 / 13.7 42.2 / 20.4 51.3 / 27.2 52.8 / 29.0 53.2 / 29.9 54.9 / 30.4 54.8 / 30.1
AnyView(30) 30.5 / 12.7 39.8 / 20.1 52.4 / 29.2 57.1 / 31.6 59.1 / 35.2 59.9 / 36.8 60.1 / 37.4
AnyView(50) 32.7 / 15.3 45.0 / 22.0 53.8 / 30.5 57.5 / 33.5 59.5 / 35.7 60.6 / 35.7 60.7 / 35.8
Continuous ImVoteNet 31.6 / 13.5 18.1 / 8.8 25.2 / 13.5 30.4 / 15.7 40.5 / 18.4 43.1 / 20.0 43.3 / 20.4
3DETR-m(10) 29.8 / 12.5 19.9 / 9.9 27.7 / 15.3 34.0 / 18.1 44.2 / 24.1 48.7 / 27.8 48.9 / 27.7
3DETR-m(30) 29.1 / 13.1 18.5 / 9.6 26.6 / 15.1 35.4 / 19.7 48.0 / 29.4 54.0 / 33.6 55.6 / 34.7
3DETR-m(50) 27.7 / 12.8 17.9 / 9.7 26.6 / 15.8 34.9 / 19.8 46.9 / 28.8 52.7 / 33.0 54.7 / 35.3
AnyView(10) 31.3 / 13.7 20.6 / 10.6 30.2 / 15.7 36.0 / 17.7 47.9 / 26.1 53.6 / 28.7 54.8 / 30.1
AnyView(30) 30.5 / 12.7 19.9 / 10.1 29.9 / 16.5 36.8 / 21.4 52.3 / 30.0 57.1 / 34.3 60.1 / 37.4
AnyView(50) 32.7 / 15.3 22.1 / 11.3 32.9 / 18.6 40.6 / 21.8 54.4 / 29.8 58.7 / 34.4 60.7 / 35.8
Figure 3: Visual results on ScanNet (IoU=0.25). We compare the predictions of 3DETR-m and AnyView after NMS with the ground-truth bounding boxes on different scales of input. Top: the whole scene consisting of 50 views, Bottom: a part of scene consisting of 10 adjacent views.